There are two fruit stalls that are located on separate areas in the park. At the end of the day, the first stall sold 25 apples and 10 oranges for a total of 610 pesos. The second stall sold 13 apples and 18 oranges for a total of 506 pesos. Find how much one orange cost
Answer by richwmiller(9135)

25a+10r=610
13a+18r=506
r=59/4= 14.75 pesos per orange
a=37/2=18.50 pesos